OpenStudy (anonymous):

Rewrite the number in decimal form: 6.15 x 10^2

OpenStudy (butterflydreamer):

hint :) 10^2 = 10 * 10 = 100 So you are working out 6.15 * 100 = ?

OpenStudy (anonymous):

this would help walpha alpha

OpenStudy (anonymous):

wouldn't that just be 615? :/ It says my answer has to be in decimal form...

OpenStudy (mahmood786123):

6.15

OpenStudy (mahmood786123):

what grade you in

OpenStudy (mahmood786123):

@EclipsedStar

OpenStudy (mahmood786123):

@kidrah69 help

OpenStudy (butterflydreamer):

well.. if you wrote your question correctly " 6.15 x 10^2" then it will just be 6.15 times 100 = 615 ... unless you're missing something? :O

OpenStudy (anonymous):

No... My question is how it's stated in the book.. Weird..

OpenStudy (butterflydreamer):

yep :) then it's just 615 xD

OpenStudy (anonymous):

ok, thanks haha :)

OpenStudy (butterflydreamer):

no problem :D!